# Active particle motion in Poiseuille flow through rectangular channels

### v1

@@ Abstract
We investigate the dynamics of a point-like active particle suspended in fluid flow through a straight channel. For this particle-fluid system, we derive a constant of motion for a general unidirectional fluid flow, and apply it to an approximation of Poiseuille flow through rectangular cross-sections. For a given rectangular cross-section, this results in a $4$D nonlinear conservative dynamical system with one constant of motion and a dimensionless parameter as the ratio of maximum flow speed to intrinsic active particle speed. We observe a diverse set of active particle trajectories with variations in system parameters and initial conditions which we classify into different types of swinging, trapping, tumbling and wandering motion. Regular (periodic/quasiperiodic) motion as well as chaotic active particle motion are observed for these trajectories and quantified using largest Lyapunov exponents. We explore the transition to chaotic motion using Poincaré maps and show “sticky” chaotic tumbling trajectories that have long transients near a periodic state. Outcomes of this work may have implications for dynamics of natural and artificial microswimmers in experimental microfluidic channels that typically have rectangular cross-sections.
@@ I Introduction
Active particles are entities that take energy from the environment and convert it into persistent motion. Examples include macroscopic living organisms, such as birds, fish and mammals, who consume energy from food and self-propel via various modes of locomotion. Active particles are also ubiquitous in the microscopic living world such as bacteria, cells, algae and other microorganisms [1]. Although persistent motion is a visible feature that is commonly associated with life, active particles also emerge in several non-equilibrium inanimate physical and chemical systems [2, 3, 4, 5, 6].
@@
Active particles immersed in a fluid medium at the micro scale, also known as microswimmers, are a commonly studied class of active particles [7]. These microswimmers ubiquitously interact with external fluid flows in various situations. For example, microswimmers routinely experience unidirectional flows in confined channels such as sperm cells swimming in fallopian tubes [8, 9], pathogens moving through blood vessels [10] and micro-robots programmed for targeted drug delivery applications [11]. In these scenarios, the coupling between external flow fields and intrinsic velocity of the active particle can lead to rich dynamical behaviors [12]. Understanding of the active particle dynamics arising from coupling with external fluid flows is not only crucial from a biological perspective, but it will also guide in design of artificial microswimmers for biomedical applications of cell manipulation, targeted drug delivery and cargo transport [13]. Further, it can guide designs of industrial and biomedical microfluidic devices aimed at focusing, sorting and filtering of microorganisms in a fluid suspension [14, 15].
@@
Zöttl and Stark [16] studied the motion of a microswimmer in unidirectional confined flows by modeling the active particle as a spherically symmetric point with constant intrinsic velocity. For $2$D planar Poiseuille flow, they showed that the equation of motion for the active particle can be mapped onto the mathematical equation of a simple pendulum, where the oscillating and circling solutions of the pendulum motion corresponds to two different types of active particle motion, swinging and tumbling, respectively. In swinging motion the upstream-oriented active particle performs oscillations about the channel centerline, whereas in tumbling motion, the active particle oscillates near the edges of the channel with fluctuating orientation and does not cross the channel centerline. Variations of this model that include additional attributes to the active particle and/or fluid flow have been investigated in detail for $2$D channel flows [17, 18, 19, 20, 21, 22]. For $3$D cylindrical Poiseuille flow, Zöttl and Stark [16] showed that the particle-fluid dynamical system is Hamiltonian with enough conserved quantities to make the system integrable. In this case, they showed that the active particle exhibits periodic motion with $3$D generalizations of swinging and tumbling trajectories. The effect of flow anisotropy was also studied by Zöttl and Stark [22] where they considered an elliptical channel cross-section and showed that the active particle motion is much more complex with typically quasiperiodic trajectories, and a few examples of chaotic motion using Poincaré maps were also reported by Zöttl [23].
@@
Although the axisymmetric fluid flow profile in a cylindrical channel results in simplified equations for the active particle motion, in many microfluidic applications concerned with natural and artificial microswimmers, microchannels with rectangular cross-sections are more commonly used since they are relatively easy to fabricate [24]. Motivated by this, in this paper, we apply the simple model of Zöttl and Stark [16] to explore the dynamics of a simple active particle in Poiseuille flow through a straight channel having square/rectangular cross-sections. The introduction of a square/rectangular cross-section introduces anisotropy by breaking the continuous rotational symmetry of fluid flow in a circular cross-section and makes the system non-integrable. We observe a rich variety of active particle motion with both quasiperiodic and chaotic trajectories that we explore in detail as a function of system parameters and initial conditions.
@@
The paper is organized as follows. In Sec. II we present the equations of motion for the particle-fluid system and derive general constants of motion for an active particle in unidirectional fluid flow. We then identify equilibrium states for an active particle in Poiseuille flow with rectangular cross-section and determine their stability in Sec. III. After briefly reviewing active particle motion in a channel with circular cross-section in Sec. IV, we present in Sec. V a detailed exploration of active particle dynamics in a channel with square cross-section. This includes a classification of trajectories, comparison with dynamics in a circular cross-section, a detailed parameter space exploration of cross-sectional active particle dynamics as well as an investigation of the transport of active particle along the channel. In Sec. VI we briefly explore the effect of aspect ratio of the rectangular cross-section on active particle motion. We provide our conclusions in Sec. VII.
@@
Figure 1: Schematic of the particle-fluid system in dimensionless units. A simple point-like active particle located at $\mathbf{r}=(x,y,z)$ and having a unit intrinsic speed in the direction of its orientation $\mathbf{\hat{e}}$ is suspended in a unidirectional channel flow $\bar{u}(x,y)$ through a straight $3$D rectangular channel with width to height aspect ratio $AR$. The particle’s orientation is represented using spherical co-ordinates with polar angle $\theta\in(-\pi/2,\pi/2)$ measuring the orientation relative to the $x-z$ plane, and azimuthal angle $\phi\in(-\pi,\pi]$ measuring angle within the $x-z$ plane relative to the (negative) $z$ axis. The left panel shows the top-view of the channel in the $x-z$ plane while the right panel shows the cross-sectional view of the channel in the $y-x$ plane.
@@ II Equations of motion
As illustrated in Fig. 1, consider a point-like active particle model of a spherical microswimmer having a constant intrinsic swimming speed $v_{0}$ in the direction of its orientation $\mathbf{\hat{e}}=e_{x}\mathbf{\hat{i}}+e_{y}\mathbf{\hat{j}}+e_{z}\mathbf{\hat{k}}$. The active particle is located at $\mathbf{r}=x\mathbf{\hat{i}}+y\mathbf{\hat{j}}+z\mathbf{\hat{k}}$ and is suspended in a steady unidirectional flow $u(x,y)\mathbf{\hat{k}}$ through a straight $3$D channel. The equations of motion for the active particle are given by [16]:
$\frac{\text{d}\mathbf{r}}{\text{d}t}=v_{0}\mathbf{\hat{e}}+u(x,y)\mathbf{\hat{k}},$ (1)
$\frac{\text{d}{\mathbf{\hat{e}}}}{\text{d}t}=\frac{1}{2}(\nabla\times u(x,y)\mathbf{\hat{k}})\times\mathbf{\hat{e}}.$ (2)
@@
Equation (1) describes the translational motion of the active particle as a combination of its intrinsic velocity $v_{0}\mathbf{\hat{e}}$ and the velocity of the background fluid flow at the particle location $u(x,y)\mathbf{\hat{k}}$, whereas Eq. (2) describes the evolution of the active particle’s orientation based on the local flow vorticity. We assume that the active particle is small compared to the cross-sectional dimensions of the channel, and hence the particle does not disturb the fluid flow. Further, we assume that the active particle stays away from bounding walls so that we can neglect interactions/collisions between the active particle and the walls.
@@
Non-dimensionalizing Eqs. (1) and (2) with a characteristic length scale $H$ of the cross-section and time scale $H/v_{0}$, we get the following dimensionless equations:
$\frac{\text{d}\mathbf{r}}{\text{d}t}=\mathbf{\hat{e}}+{\bar{u}(x,y)\mathbf{\hat{k}}},$
$\frac{\text{d}{\mathbf{\hat{e}}}}{\text{d}t}=\frac{1}{2}(\nabla\times{\bar{u}}(x,y)\mathbf{\hat{k}})\times\mathbf{\hat{e}}.$
@@
Here, ${\bar{u}}(x,y)$ is the dimensionless flow field (scaled by the active particle speed $v_{0}$). In Cartesian components, we get six nonlinear ordinary differential equations (ODEs) as follows:
$\dot{x}=e_{x},\,\dot{y}=e_{y},\,\dot{z}=e_{z}+\bar{u}(x,y),$ (3)
$\dot{e}_{x}=-\frac{1}{2}e_{z}\frac{\partial\bar{u}}{\partial x},\,\dot{e}_{y}=-\frac{1}{2}e_{z}\frac{\partial\bar{u}}{\partial y},\,\dot{e}_{z}=\frac{1}{2}e_{x}\frac{\partial\bar{u}}{\partial x}+\frac{1}{2}e_{y}\frac{\partial\bar{u}}{\partial y}.$
@@
We note from Eq. (3) that the dynamical flow is divergence free, that is $\nabla\cdot(\dot{x},\dot{y},\dot{z},\dot{e}_{x},\dot{e}_{y},\dot{e}_{z})=0$. Hence, the dynamical system is conservative and phase-space volumes are preserved under the dynamical flow. We further note that the $z$ variable can be decoupled i.e. the $\dot{z}$ equation can be integrated separately, thus reducing our dynamical system to five equations. The effective dimension of our dynamical system is further reduced by identifying constants of motion, i.e. quantities that remain constant during the evolution of the system. A trivial constant of motion for our system is
@@
since the orientation vector has unit magnitude. We have also identified another constant of motion as (see Appendix A for a proof)
$H_{g}=-\frac{1}{2}\bar{u}(x,y)+e_{z}.$ (5)
With these two constants of motion, our five-dimensional dynamical system reduces to three effective dimensions. We can explicitly use the constant of motion in Eq. (4) and reduce our system to four nonlinear ODEs. This is done by parameterizing the three Cartesian orientation components of the active particle, $e_{x},e_{y}$ and $e_{z}$ using spherical co-ordinate angles $\theta$ and $\phi$ as follows:
@@
$\dot{x}$ $=-\cos\theta\sin\phi,$ (6)
$\dot{y}$ $=\sin\theta,$
$\dot{\theta}$ $=\frac{1}{2}\frac{\partial\bar{u}}{\partial y}\cos\phi,$
$\dot{\phi}$ $=\frac{1}{2}\frac{\partial\bar{u}}{\partial y}\tan\theta\,\sin\phi-\frac{1}{2}\frac{\partial\bar{u}}{\partial x},$
along with one constant of motion in Eq. (5) which is rewritten in the above parameterization as
$H_{g}=-\frac{1}{2}\bar{u}(x,y)-\cos\phi\cos\theta.$ (7)
We note that up to this point, our consideration of the fluid flow field $\bar{u}(x,y)$ has been general and hence the constant of motion in Eq. (5) exists independent of the specific flow profile.
We now consider the specific fluid flow profile of Poiseuille flow in a $3$D straight channel having a rectangular cross-section with width $W$ and height $H$. We approximate this flow profile by
${u}(x,y)=U_{0}\left(1-\left(\frac{x}{W}\right)^{2}\right)\left(1-\left(\frac{y}{H}\right)^{2}\right).$
This expression provides a good approximation to the exact solution of Poiseuille flow in rectangular channels expressed as an infinite series (see Appendix B). In dimensionless form (with channel height $H$ as the length scale and $H/v_{0}$ as the time scale), our flow field is given by
${\bar{u}}(x,y)=\bar{U}\left(1-\left({\frac{x}{AR}}\right)^{2}\right)\left(1-y^{2}\right),$ (8)
where $\bar{U}=U_{0}/v_{0}$, is the ratio of maximum flow speed to the intrinsic active particle speed and $AR=W/H$, is the aspect ratio of the rectangular cross-section. Substituting this flow field in Eq. (6) we get the following $4$D dynamical system (along with the constant of motion in Eq. (7)):
$\dot{x}$ $=-\cos\theta\sin\phi,$ (9)
$\dot{y}$ $=\sin\theta,$
$\dot{\theta}$ $=-\bar{U}y\cos\phi\left(1-\left({\frac{x}{AR}}\right)^{2}\right),$
$\dot{\phi}$ $=-\bar{U}y\tan\theta\sin\phi\left(1-\left({\frac{x}{AR}}\right)^{2}\right)+\bar{U}\frac{x}{AR^{2}}(1-y^{2}),$
where $-AR<x<AR$, $-1<y<1$, $-\pi/2<\theta<\pi/2$, $-\pi<\phi\leq\pi$ and $z\in\mathbb{R}$.

### v2

@@ Abstract
We investigate the dynamics of a point-like active particle suspended in fluid flow through a straight channel. For this particle-fluid system, we derive a constant of motion for a general unidirectional fluid flow, and apply it to an approximation of Poiseuille flow through channels with rectangular cross-sections. We obtain a $4$D nonlinear conservative dynamical system with one constant of motion and a dimensionless parameter describing the ratio of maximum flow speed to intrinsic active particle speed. Applied to square channels, we observe a diverse set of active particle trajectories with variations in system parameters and initial conditions which we classify into different types of swinging, trapping, tumbling and wandering motion. Regular (periodic/quasiperiodic) motion as well as chaotic active particle motion are observed for these trajectories and quantified using largest Lyapunov exponents. We explore the transition to chaotic motion using Poincaré maps and show “sticky” chaotic tumbling trajectories that have long transients near a periodic state. We briefly illustrate how these results extend to rectangular cross-sections with width/height ratio larger than one. Outcomes of this work may have implications for dynamics of natural and artificial microswimmers in experimental microfluidic channels that typically have rectangular cross-sections.
@@ I Introduction
Active particles are entities that take energy from the environment and convert it into persistent motion. Examples include macroscopic living organisms, such as birds, fish and mammals, which consume energy from food and self-propel via various modes of locomotion. Active particles are also ubiquitous in the microscopic living world such as bacteria, cells, algae and other microorganisms [1]. Although persistent motion is a visible feature that is commonly associated with life, active particles also emerge in several non-equilibrium inanimate physical and chemical systems [2, 3, 4, 5, 6].
@@
Active particles immersed in a fluid medium at the micro scale, also known as microswimmers, are a commonly studied class of active particles [7]. These microswimmers ubiquitously interact with external fluid flows in various situations. For example, microswimmers routinely experience unidirectional flows in confined channels such as sperm cells swimming in fallopian tubes [8, 9], pathogens moving through blood vessels [10] and micro-robots programmed for targeted drug delivery applications [11]. In these scenarios, the coupling between external flow fields and intrinsic velocity of the active particle can lead to rich dynamical behaviors [12]. Understanding of the active particle dynamics arising from coupling with external fluid flows is not only interesting from a biological perspective, but is also crucial for design of artificial microswimmers for biomedical applications of cell manipulation, targeted drug delivery and cargo transport [13]. Further, it can aid design of industrial and biomedical microfluidic devices aimed at focusing, sorting and filtering of microorganisms in a fluid suspension [14, 15].
@@
Zöttl and Stark [16] studied the motion of a microswimmer in unidirectional confined flows by modeling the active particle as a spherically symmetric point with constant intrinsic velocity. For $2$D planar Poiseuille flow, they showed that the equation of motion for the active particle can be mapped onto the mathematical equation of a simple pendulum, where the oscillating and circling solutions of the pendulum motion correspond to two different types of active particle motion, swinging and tumbling, respectively. In swinging motion the upstream-oriented active particle performs oscillations about the channel centerline, whereas in tumbling motion, the active particle oscillates near the edges of the channel with fluctuating orientation and does not cross the channel centerline. Variations of this model that include additional attributes to the active particle and/or fluid flow have been investigated in detail for $2$D channel flows [17, 18, 19, 20, 21, 22]. For $3$D cylindrical Poiseuille flow, Zöttl and Stark [16] showed that the particle-fluid dynamical system is Hamiltonian with enough conserved quantities to make the system integrable. In this case, they showed that the active particle exhibits periodic motion with $3$D generalizations of swinging and tumbling trajectories. The effect of flow anisotropy was also studied by Zöttl and Stark [22] who showed, for an elliptical channel cross-section, that the active particle motion is much more complex with, typically, quasiperiodic trajectories. Using Poincaré maps, a few examples of chaotic motion were also reported by Zöttl [23].
@@
Although the axisymmetric fluid flow profile in a cylindrical channel results in simplified equations for the active particle motion, in many microfluidic applications concerned with natural and artificial microswimmers, microchannels with rectangular cross-sections are more commonly used since they are relatively easy to fabricate [24]. Motivated by this, herein we apply the model of Zöttl and Stark [16] to explore the dynamics of a simple active particle suspended in Poiseuille flow through a straight channel having square/rectangular cross-section. The introduction of a square/rectangular cross-section introduces anisotropy, by breaking the continuous rotational symmetry of fluid flow that exists in a circular cross-section, and makes the system non-integrable. We observe a rich variety of active particle motion with both quasiperiodic and chaotic trajectories. These motions are explored in detail as a function of system parameters and initial conditions.
@@
The paper is organized as follows. In Sec. II we present the equations of motion for the particle-fluid system and derive general constants of motion for an active particle in unidirectional fluid flow. We then, in Sec. III, identify equilibrium states for an active particle suspended in Poiseuille-like flow through a rectangular cross-section and determine their stability. After briefly reviewing the special case of active particle motion in a channel with circular cross-section in Sec. IV, we present a detailed exploration of active particle dynamics in a channel with square cross-section in Sec. V. This includes a classification of trajectories, comparison with dynamics in a circular cross-section, a detailed parameter space exploration of cross-sectional active particle dynamics, as well as an investigation of the transport of an active particle along the channel. In Sec. VI we briefly explore the effect of the width/height ratio of the rectangular cross-section on active particle motion. We provide our conclusions in Sec. VII.
@@
Figure 1: Schematic of the particle-fluid system. A simple point-like active particle located at $\mathbf{r}=(x,y,z)$ and having a constant intrinsic speed in the direction of its orientation $\mathbf{\hat{e}}$ is suspended in a unidirectional channel flow ${u}(x,y)$ through a straight $3$D rectangular channel with width to height aspect ratio $AR$. The particle’s orientation is represented using spherical co-ordinates with polar angle $\theta\in(-\pi/2,\pi/2)$ measuring the orientation relative to the $x-z$ plane, and azimuthal angle $\phi\in(-\pi,\pi]$ measuring angle within the $x-z$ plane relative to the (negative) $z$ axis. The left panel shows the top-view of the channel in the $x-z$ plane while the right panel shows the cross-sectional view of the channel in the $y-x$ plane.
@@ II Equations of motion
Consider the point-like active particle model of a spherical microswimmer illustrated in Fig. 1. The active particle has a constant intrinsic swimming speed $v_{0}$ in the direction of its orientation $\mathbf{\hat{e}}=e_{x}\mathbf{\hat{i}}+e_{y}\mathbf{\hat{j}}+e_{z}\mathbf{\hat{k}}$, is located at $\mathbf{r}=x\mathbf{\hat{i}}+y\mathbf{\hat{j}}+z\mathbf{\hat{k}}$, and is suspended in a steady unidirectional flow $u(x,y)\mathbf{\hat{k}}$ through a straight $3$D channel. The equations of motion for the active particle are given by [16]:
$\frac{\text{d}\mathbf{r}}{\text{d}t}$ $=v_{0}\mathbf{\hat{e}}+u(x,y)\mathbf{\hat{k}},$ (1a)
$\frac{\text{d}{\mathbf{\hat{e}}}}{\text{d}t}$ $=\frac{1}{2}(\nabla\times u(x,y)\mathbf{\hat{k}})\times\mathbf{\hat{e}}.$ (1b)
@@
Equation (1a) describes the translational motion of the active particle as a combination of its intrinsic velocity $v_{0}\mathbf{\hat{e}}$ and the local velocity of the background fluid flow $u(x,y)\mathbf{\hat{k}}$, whereas Eq. (1b) describes the evolution of the active particle’s orientation based on the local flow vorticity. We assume that the active particle is small compared to the cross-sectional dimensions of the channel, and hence the particle does not disturb the fluid flow. Further, we assume that the active particle stays away from bounding walls so that we can neglect interactions/collisions between the active particle and the walls.
@@
Non-dimensionalizing Eqs. (1a) and (1b) with a characteristic length scale $H$ of the cross-section and time scale $H/v_{0}$, we obtain the following dimensionless equations:
$\frac{\text{d}\bar{\mathbf{r}}}{\text{d}\bar{t}}$ $=\mathbf{\hat{e}}+{\bar{u}(\bar{x},\bar{y})\mathbf{\hat{k}}},$ (2a)
$\frac{\text{d}{\mathbf{\hat{e}}}}{\text{d}\bar{t}}$ $=\frac{1}{2}(\bar{\nabla}\times{\bar{u}}(\bar{x},\bar{y})\mathbf{\hat{k}})\times\mathbf{\hat{e}}.$ (2b)
@@
Here, the dimensionless variables are denoted with an overbar and the dimensionless flow field ${\bar{u}}(\bar{x},\bar{y})$ is scaled with the active particle speed $v_{0}$. We now drop the overbars on dimensionless variables for convenience. In component form, we get a system of six nonlinear ordinary differential equations (ODEs) as follows:
$\dot{x}$ $=e_{x},$ $\dot{e}_{x}$ $=-\frac{1}{2}e_{z}\frac{\partial{u}}{\partial x},$
$\dot{y}$ $=e_{y},$ $\dot{e}_{y}$ $=-\frac{1}{2}e_{z}\frac{\partial{u}}{\partial y},$
$\dot{z}$ $=e_{z}+{u}(x,y),$ $\dot{e}_{z}$ $=\frac{1}{2}e_{x}\frac{\partial{u}}{\partial x}+\frac{1}{2}e_{y}\frac{\partial{u}}{\partial y}.$
We note from the component form that the dynamical flow is divergence free, that is
$\frac{\partial\dot{x}}{\partial x}+\frac{\partial\dot{y}}{\partial y}+\frac{\partial\dot{z}}{\partial z}+\frac{\partial\dot{e}_{x}}{\partial e_{x}}+\frac{\partial\dot{e}_{y}}{\partial e_{y}}+\frac{\partial\dot{e}_{z}}{\partial e_{z}}=0.$
@@
Hence, the dynamical system is conservative and phase-space volumes are preserved under the dynamical flow. We further note that the $z$ variable can be decoupled, i.e. the $\dot{z}$ equation can be integrated separately, thus reducing our dynamical system to five differential equations. The effective dimension of our dynamical system is further reduced by identifying constants of motion, i.e. quantities that remain constant during the evolution of the system. A trivial constant of motion for our system is
@@
since the orientation vector maintains unit magnitude. We have also identified a second constant of motion as (see Appendix A for a proof)
$H_{g}=-\frac{1}{2}{u}(x,y)+e_{z}.$ (4)
With these two constants of motion, our five-dimensional dynamical system reduces to three effective dimensions. We can implicitly use the constant of motion in Eq. (3) and reduce our system to four nonlinear ODEs by parameterizing the Euler axis using spherical co-ordinate angles $\theta$ and $\phi$ as follows:
@@
$\dot{x}$ $=-\cos\theta\sin\phi,$ (5a)
$\dot{y}$ $=\sin\theta,$ (5b)
$\dot{\theta}$ $=\frac{1}{2}\frac{\partial{u}}{\partial y}\cos\phi,$ (5c)
$\dot{\phi}$ $=\frac{1}{2}\frac{\partial{u}}{\partial y}\tan\theta\,\sin\phi-\frac{1}{2}\frac{\partial{u}}{\partial x},$ (5d)
along with the constant of motion in Eq. (4) which is rewritten in the above parameterization as
$H_{g}=-\frac{1}{2}{u}(x,y)-\cos\phi\cos\theta.$ (6)
We note that up to this point, our consideration of the fluid flow field ${u}(x,y)$ has been general and hence the constant of motion in Eq. (4) exists independent of the specific flow profile.
We now consider the specific fluid flow profile of Poiseuille flow in a $3$D straight channel having a rectangular cross-section with width $W$ and height $H$. With $H$ as the length scale and defining the aspect ratio $AR=W/H$, we approximate the dimensionless flow profile by
${u}(x,y)=U\left(1-\left(\frac{x}{AR}\right)^{2}\right)\left(1-y^{2}\right),$ (7)
where $U=u(0,0)$ is the maximum velocity in the channel scaled with the intrinsic particle speed $v_{0}$. This expression provides a good approximation to the exact solution of Poiseuille flow in rectangular channels expressed as an infinite series (see Appendix B). Substituting this flow field in Eq. (5) we get the following $4$D dynamical system (along with the constant of motion in Eq. (6)):
$\dot{x}$ $=-\cos\theta\sin\phi,$ (8a)
$\dot{y}$ $=\sin\theta,$ (8b)
$\dot{\theta}$ $=-{U}y\cos\phi\left(1-\frac{x^{2}}{AR^{2}}\right),$ (8c)
$\dot{\phi}$ $=-{U}y\tan\theta\sin\phi\left(1-\frac{x^{2}}{AR^{2}}\right)+{U}\frac{x(1-y^{2})}{AR^{2}},$ (8d)
where $-AR<x<AR$, $-1<y<1$, $-\pi/2<\theta<\pi/2$ and $-\pi<\phi\leq\pi$.